Let 𝑓 be differentiable. Show that if lim𝑥→∞ 𝑓(𝑥) = 𝐿 ∈ ℝ then lim𝑥→∞ 𝑓 ′ (𝑥) = 0. Provided that the latter limit is existing. Give an example where the converse is not true. Also give an example for which the limit of 𝑓 ′ is not existing even though the limit of 𝑓 is the same as given.


1
Expert's answer
2021-10-08T15:49:05-0400

Let "f\in L^1(R),\exist\space lim_{x->\infty}f(x)=b\neq0"

Let "\epsilon\space be\space choosen\space as\space0<\epsilon<|b|/2"

According with definition of lim

"\exist M>0:\forall(x>M)|f(x)-b|<\epsilon<|b|/2"

We may think b>0 taking otherwise -f(x)

Then f(x)>b/2 as x>M and "\int\limits_{M}^{M\cdot K} |f(x)| dx>\frac{M\cdot K\cdot b}{2}"

with any K>1.

Also valid that "\int\limits_{-\infty}^{\infty} |f(x)| dx\ge 
\int\limits_{M}^{M\cdot K} |f(x)| dx>\frac{M\cdot K\cdot b}{2}"

If we will take K rather big we will have

"\int\limits_{-\infty}^{\infty} |f(x)| dx=\infty"  that is controversy


An example "f\in L_1(R)" such that "\bar{\exist}\space lim_{x->\infty}f(x)" :

We take  "f(x)=1\space if \space|k-x|\le\frac{1}{k^2},k=1,2,..." and f(x)=0 otherwise. Then "\int\limits_{-\infty}^{\infty} |f(x)| dx=""2\cdot \displaystyle\sum_{i=1}^{\infty}\frac{1}{k^2}<\infty"

but "\exists \{x_n\}_{n=1,2,..}" such that "f(x_n)=1" if n is odd and "f(x_n)=0" if n is even there fore limit f(x) on "\infty" doesn't exist.


if lim𝑥→∞ 𝑓(𝑥) = 𝐿 ∈ ℝ then:

lim𝑥→∞ 𝑓'(𝑥) =lim𝑥→∞ (dL/dx)=0